OpenStudy (anonymous):

integrates 1/x(logx)powerof n w.r.t.x

OpenStudy (anonymous):

u=ln(x)

OpenStudy (anonymous):

\[\int\limits \frac{1}{ x [\ln(x)]^n} dx\]

OpenStudy (anonymous):

if its that^

OpenStudy (anonymous):

if its that^

OpenStudy (anonymous):

let u=ln(x) du = (1/x) dx

OpenStudy (anonymous):

\[= \int\limits \frac{1}{u^n} du \]

OpenStudy (anonymous):

\[= \int\limits u^{-n} du = \frac{u^{-n+1} }{1-n} +C \]

OpenStudy (anonymous):

but that is only for \[n \neq \]

OpenStudy (anonymous):

\[= \frac{ (\ln(x))^{1-n} }{ 1-n} +C \] for \[n \neq 1\] = \[\ln(\ln(x)) +C \] for n=1